# Decoherence of electron spin qubit during transfer between two semiconductor quantum dots at low magnetic fields

### v2

@@ V Discussion and summary
where $Q^{(n)}$ is the error accumulated during transition between $(n-1)$th and $n$th dot. In the present context, single charge-transfer error even means that the electron gets left behind in one of the dots when it should transfer to the next one. However, note that in the simplest version of a scalable shuttling scheme, in which a few global voltages are used to simultaneously trigger the appropriate swings of interdot detunings in a long chain of QDs [7], an electron that got left behind will start traveling in the opposite direction in subsequent driving cycles. In a more complicated scheme it would resume its motion in the correct direction along the QD chain. However, even in the latter scenario, the delay associated with staying in a random dot (with a random deviation of Zeeman splitting from the acerage value) for one period of driving, will lead to appreciable dephasing if the delayed qubit. For simplicity, we assume here that every charge transfer failure means an irreversible loss of coherence of the shuttled qubit.
For a QD chain with uniform parameters, for which each pair of adjacent QDs has $2t_{c}^{(n)}\approx 2t_{c}^{(0)}$, the overall transfer error can be estimated as $Q_{N}=NQ^{(1)}$. Based on the result from Fig. 5, the largest considered $2t_{c}=100\mu$eV would allow for coherent transfer, defined as $Q^{(N)}<10^{-3}$, across $N=1000$ and $N=10-100$ dots in Si and GaAs ($|a|=1\,\mu$eV) respectively. With a distance between the dots of $d=100$nm this corresponds to $L=100\mu$m and $L=1-10\mu$m respectively, which is consistent with requirement posed by [2]. From our results for smaller $2t_{c}$, it is clear that $2t_{c}=40\mu$eV would not allow for a transfer across $L\!=\!10\mu$m even in a perfectly uniform QD chain. We predict that the minimal tunnel coupling required for such a transfer is $2t_{c}\approx 60\mu$eV, for which high precision in choosing $v$ is required, while the resulting fidelity depends on the details of experimental device, i.e. temperature, $T_{2}^{*}$, and the amplitude and spatial correlations of charge noise.
In a more realistic situation of QD chain with tunnel couplings distributed with appreciable standard deviation around a mean value that is “large enough” for low coherent transfer error, the transfer can be limited by a few weak links, or even dominated by a single weakly coupled pair, i.e. $Q_{N}\approx\text{max}(Q)$. Again, based on results from Fig. 5 this puts constrains on the weakest tunnel coupling along the chain to $2t_{c}\approx 40\,\mu$eV as achievable error was slightly above $Q^{(1)}=10^{-3}$ in GaAs and slightly below in Si. In such a case, the presence of second minimum at low $v$ in Si can be in principle used to improve the transfer error by adjusting the sweep rate for the whole transfer, or only when the qubit is passing through the weak link. However the latter would pose a challenge for the electronic control, and also require a characterization of the chain, while the former would significantly slow down the electron transfer, exposing the qubit to stronger dephasing due to local low-frequency spin splitting fluctuations, i.e. errors described by Eq. (30).
The additivity of errors might no longer hold for dephasing caused by spatiotemporally correlated low-frequency noise, due to possibly coherent sum of random phases:
@@
which contribution to an error after averaging depends on the correlations $\langle\delta\varphi^{2}\rangle=\sum_{nm}\langle\delta\varphi^{(n)}\delta\varphi^{(m)}\rangle$. However for a linear chain, each dot is visited only once, which together with only weak correlation of the noise in nearest neighbours $\langle\delta\varphi^{(n)}\delta\varphi^{(n+1)}\rangle\ll\langle[\delta\varphi^{(n)}]^{2}\rangle$ is expected to introduce a small modification to linear scaling, such that for uniform case of $\langle[\delta\varphi^{(n)}]^{2}\rangle\approx Q_{\varphi,1}$ we can estimate
@@
where we introduced constant numerical factor $\xi\gtrsim 1$ encapsulating spatial correlation of charge noise between nearest and next-nearest neighbours [55]. The above is the manifestation of motional narrowing process, which in comparison to stationary system, decreases coherent error by the virtue of interacting with many uncorrelated environments [57]. Such motional narrowing leading to decreased dephasing of moving qubit compared to the stationary one was observed in recent shuttling experiment with a moving QD in Si/SiGe [9].
@@
In view of latest experimental results in Ge holes [15], we highlight that correlations of low-frequency noise are expected to be more relevant for sequential transition between few quantum dots. Since in [15] tunnel coupling $2t_{c}\approx 100\mu$eV, allowed for $v\approx 10^{3}\,\mu$eV/ns, no charge excitations are expected there, and the error should come from dephasing noise caused by very large $\Delta E_{Z}$ specific to holes in Ge [58, 59, 60]. However, in such a setup of shuttling back and forth between two or three dots, the same dot is visited multiple times, meaning that even without spatial correlations of the noise, slow fluctuations $\delta\varphi$ can add coherently, which can result in $\langle\delta\varphi^{2}\rangle>N^{\alpha}$, with $\alpha>1$, scaling. Moreover, repetition of a shuttling process could create a noise filter that commensurate with shuttling frequency. Thus, predicting exact relation between spectral density of the noise, electron path and dephasing would require a separate, careful analysis outside of the scope of this paper.
In summary, we have analyzed possibility of coherent electron transfer between two tunnel-coupled semiconductor quantum dots. We analyzed contrasting examples of spin qubits realised in GaAs and Si nanostructures. We have shown that maintaining qubit coherence would benefit from using large detuning sweep rates, that according to the Landau-Zener formula are enabled by maximising the tunnel coupling. Slower sweeps activates a number of error mechanisms, leading to charge excitation, spin relaxation, and spin dephasing. In the process we have identified the origins of such errors in a form of inelastic transition between the instantaneous levels, slow fluctuations of detuning both activated by the dot-dependent Zeeman splitting, and spin relaxation induced by the spin-orbit interaction. To provide qualitative result we have modelled environment for the relevant case of Si and GaAs quantum dots. In general, in both devices an optimal sweep rate was found, as a result of the non-monotonic dependence of $Q$ on sweep rate $v$. Typically, the trade-off resulting in a local minimum of $Q(v)$ is between the increasing probability of losing the electron charge at high $v$, and spin relaxation or spin dephasing, becoming more dangerous at low $v$.
In GaAs, longer transfer time expose the electron spin to the inhomogenous dephasing associated with the $T_{2}^{*}$. In principle this issue can be fixed when $T_{2}^{*}$ is enhanced, for instance by continuous estimation of the nuclear field [41, 42, 61]. However, as we have shown in this paper, this improvement is limited by another source of error: the spin-relaxation due to non-negligible spin-orbit coupling, which dominates the error at lower $v$. We have speculated that even for optimized DQD geometry, i.e. weaker spin-orbit interaction, the error is unlikely to allow for coherent transfer at $\mu$m distance.
In contrast, we have found that in Si-based devices the transfer error as a function of $v$ exhibits more interesting structure, with up to two local minima. The first of them has been already identified with the interplay between the coherent and incoherent charge excitation, that result in leaving the electron in the initial dot [23, 25]. Here we have found an increase of the phase error due to charge noise activated in the presence of interdot difference in Zeeman splittings, which has to be taken into account in case of spin shuttling in vicinity of the manipulation region. We have reported that away from this region, where the values of the spin splitting difference are not enhanced, similar contribution is shifted to lower $v$, at which its influence becomes comparable to that of low-frequency fluctuations of the local spin splitting. For Si, this takes place at low sweep rates, at which the charge transfer becomes inelastic and decreasing with $v$. As a result a second minimum of an error at much lower $v$ can be created, however as we have shown no significant improvement in comparison to the other minimum is expected.
Concluding, we have analyzed the most relevant mechanism leading to error in coherent electron spin qubit transfer between two Si- and GaAs- quantum dots. We have found a trade-off behaviour between the mechanisms affecting charge transfer and electron spin phase, as typically the latter one is lost for the slowest sweeps. We have identified local minima of the error in coherent transfer, and analyzed their implication for the coherent transfer across few $\mu$m. We have proven that coherent shuttling using pre-defined quantum dots in isotopically purified Si is more promising in comparison to GaAs, even in presence of nuclear field narrowing in the latter. We have found that in silicon, coherent transfer across few $\mu$m would require $2t_{c}>40\,\mu$eV at the weakest link and $2t_{c}>60\,\mu$eV for majority of dots pair. This requirement for hundred-dot array makes this method of transfer challenging for future designs.
@@ Appendix B Relaxation rates
where $S_{i}(\Omega)$ is the spectral density of the environment, while the time-dependence of the relaxation rate $\Gamma_{sq\to s^{\prime}q^{\prime}}(t)$, that results from the time-depedence of the states |$\ket{\psi_{sq}(t)}$ and the energies $E_{sq}(t)$ was omitted for brevity.
@@
Using definition of the instatnanous states Eq. (LABEL:), we have:
@@ Appendix C Relaxation rates in Si and GaAs: Spectral densities
where we use $Z=50\Omega$ and $R_{q}=\pi/e^{2}\approx 13k\Omega$.
@@
Figure 8: Spin relaxation rate in GaAs (a) and Si (b). In both panels we compare two tunnel couplings ($2t_{c}=40\mu$eV (solid line), $2t_{c}=100\mu$eV (dashed), as well as two spin splittings $E_{Z}=30\mu$eV (line without dots), $E_{Z}=10\mu$eV (lines with the dots). Finally for GaAs (a) we compare contributions from intrinsic SOC, i.e $|a|=1\mu$eV, $b=0$ (red) against synthetic $|a|=0,b=1\mu$eV (black). Similarly for Si (b) we compare intrinsic SOC $|a|=0.1\mu$eV, $b=0$ (green) against sythetic $|a|=0,b=1\mu$eV (yellow).
@@ Appendix D Dephasing due to low-frequency noise
D.1 Numerical test
In this section we perform numerical test of an adiabatic dephasing. To do so we average $\varphi$, given by Eq. (87), over $\delta\epsilon$, $\delta t_{c}$, $\delta E_{Z,L}$, $\delta E_{Z,R}$. The results are depicted in Fig. 9, where numerically simulated $\langle\delta\varphi^{2}\rangle$ (dots) is shown as a function of noiseless difference in spin splittings $\Delta E_{Z}$ between the two dots.

### v3

@@ V Implications for transfer of a qubit through a chain of dots
where $Q^{(n)}$ represents the error accumulated during the transition between the $(n-1)$th and $n$th dot. A single charge-transfer error implies that the electron is left behind in the initial dot when it should have transferred to the next one. However, in the simplest version of a scalable shuttling scheme — in which a few global voltages are used to simultaneously trigger the appropriate interdot detuning swings in a long chain of quantum dots (QDs) [7]—an electron left behind will begin traveling in the opposite direction in subsequent driving cycles. In a more complex scheme, it would resume moving in the correct direction along the QD chain. Nevertheless, even in this latter scenario, the delay caused by staying in a random dot (with a random deviation of Zeeman splitting from the average value) for one period of driving leads to significant dephasing of the delayed qubit. For simplicity, we assume here that each charge transfer failure results in an irreversible loss of coherence for the shuttled qubit.
For a QD chain with uniform parameters, for which each pair of adjacent QDs has $2t_{c}^{(n)}\approx 2t_{c}^{(0)}$, the overall transfer error is estimated as $Q_{N}=NQ_{1}$, where $Q_{1}$ is the error through typical transition away from manipulation region. With these assumptions, the coherent transfer of $Q_{N}<10^{-3}$ through $N=100$ dots corresponding to approximately shuttling distance of $L=10$ $\mu$m [2], would require $602t_{c}>$ 60 $\mu$eV in Si and $1002t_{c}>$ 100 $\mu$eV in GaAs, as illustrated by the lower dashed line in Fig. 6. From our results it is clear that tunnel coupling below that value, would not allow for a transfer across $L\!=\!10$ $\mu$m even in a perfectly uniform QD chain.
Due to nature of the errors, the corresponding optimal sweep rate for the coherent transfer aligns with the one that minimises the probability of excitation to the higher energy state, making the tunnel coupling requirements across the array consistent with those outlined in [31]. In that study, in absence of spin degree of freedom, it was suggested that slower sweep rates could enhance charge transfer. However, here we have shown that this strategy is unlikely to be beneficial for coherent transfer, as lowering $v$ below the value optimal for successful charge transfer triggers significant spin dephasing and relaxation.
On the other hand, when tunnel couplings between neighboring dots in a long chain exhibit significant spread around their mean value (assumed now to be larger than the above-discussed limit of $602t_{c}\!=\!$ 60 $\mu$eV) the coherent transfer error can be limited by a few weak links (dots coupled by smallest $t_{c}$, or dots in manipulation region coupled with not-large-enough $t_{c}$), or in an extreme case, by a single transition with largest error $Q$, i.e. $Q_{N}\approx\text{max}(Q)$. Based on the results from Fig. 6 this puts constrains on the weakest tunnel coupling along the chain to $2t_{c}\!>\!50\,\mu$eV in GaAs away from manipulation region (red line), which is needed for $Q\!>\!10^{-3}$. In manipulation regions, where gradients and the $g$-factor difference are likely to be significantly larger, the minimal tunnel coupling is higher and equal to $2t_{c}\!>\!60\,\mu$eV in GaAs and $2t_{c}\!>\!35\,\mu$eV in Si, as illustrated in Fig. 6 by the intersection of the upper dashed line with the red and yellow lines respectively.
While for assumed magnetic field of $E_{Z}=30\mu$eV, the optimal $Q$ was limited by the charge and relaxation errors, for smaller magnetic field a second minimum created at much lower sweep rate can give a lower $Q$ (see a case of two minima in Fig. 5 b). If the left minimum is lower, the value of lowest $Q$ becomes dominated by the dephasing caused by spatiotemporally correlated low-frequency noise, for which additivity might no longer hold. In particular, assuming dephasing is caused by the coherent sum of random phases:
@@
contribution to an error after averaging depends on the correlations $\langle\delta\varphi^{2}\rangle=\sum_{nm}\langle\delta\varphi^{(n)}\delta\varphi^{(m)}\rangle$. For a linear chain, each dot is visited only once, which is expected to introduce only a small modification to linear scaling due to only weak correlation of the noise in nearest neighbours $\langle\delta\varphi^{(n)}\delta\varphi^{(n+1)}\rangle\ll\langle[\delta\varphi^{(n)}]^{2}\rangle$. Assuming the uniform chain of dots $\langle[\delta\varphi^{(n)}]^{2}\rangle\approx Q_{\varphi,1}$ we can estimate
@@
where numerical factor $\xi\gtrsim 1$ encapsulates spatial correlation of charge noise between nearest and next-nearest neighbours [69]. The above is the manifestation of motional narrowing process, which in comparison to stationary system, decreases coherent error by the virtue of interacting with many uncorrelated environments [71]. Such motional narrowing leading to decreased dephasing of moving qubit compared to the stationary one was observed in recent shuttling experiment with a moving QD in Si/SiGe [9].
@@
In light of recent experimental results in Ge hole systems [16], we emphasize that correlations of low-frequency noise are expected to be more relevant for sequential transitions between a few quantum dots. For instance in Ref. [16] the tunnel coupling $2t_{c}\approx 100\,\mu\text{eV}$ and optimised sweep rate should prevent significant charge excitations. Instead, we expect the error to stem from relaxation or, more likely, dephasing caused by the large $\Delta E_{Z}$ specific to holes in Ge [72, 73, 74]. However, in a setup where the shuttling occurs back and forth between two or three dots, the same dot is visited multiple times. This means that, even without spatial correlations in the noise, slow fluctuations $\delta\varphi$ can accumulate coherently, potentially resulting in $\langle\delta\varphi^{2}\rangle>N^{\alpha}$ scaling, with $21<\alpha<$ 2. Moreover, repeating the shuttling process could create a noise filter that is commensurate with the shuttling frequency. Thus, predicting the exact relationship between the spectral density of the noise, the electron path, and the resulting dephasing requires a separate, detailed analysis, which lies outside the scope of this paper.
VI Summary
In conclusion, we have analyzed the most relevant mechanisms leading to errors in coherent electron spin qubit transfer between tunnel-coupled quantum dots in Si and GaAs. Although charge transfer error in the presence of realistic environmental noise can be made arbitrarily small by using small enough sweep rates $v$ [31], the same does not hold for the coherent transfer error, $Q$, to which spin relaxation and dephasing also contribute. This happens because all the discussed here mechanisms of spin dephasing and relaxation become more efficient with decreasing $v$. Channels for spin dephasing associated with interdot qubit transfer are activated in presence of nonzero difference in Zeeman splitting, $\Delta E_{Z}$, in the two dots (due to either a $g$-factor difference, or presence of magnetic field gradient) and randomness in fractions of time spent in each of the dots caused by interaction of electron charge with high- and low-frequency electric field noise. The high-frequency noise can excite the electron out of its instantaneous ground state when the energy gap between this state and the excited one is minimal - i.e. at the tunnel-coupling induced anticrossing. Even if the electron eventually relaxes due to phonon emission to the correct final state in the target dot (so that the charge transfer error is much lower that the probability of the excitation at the anticrossing), the stochastic character of this relaxation leads to a random contribution to spin phase, and thus to spin qubit dephasing. Spin relaxation is also enhanced compared to the relaxation in one of the dots when the electron is delocalized between the dots, and the mixing of ground and excited state by spin-orbit interaction (either intrinsic, or extrinsic, due to transverse component of the mangetic fild gradient) is the strongest due to the energy gap between these states being the lowest. Finally, the low-frequency noise in detuning gives a random contribution to phase of the qubit even when charge dynamics is perfectly adiabatic, and averaging over this noise give phase error that scales as $\Delta E_{Z}^{2}/v^{2}$. Clearly, large $\Delta E_{Z}$, and small tunnel coupling $t_{c}$ (which, apart from enhancing excitation probability and spin relaxation at the anticrossing, necessitates smaller $v$ to avoid nonadiabaticity of charge dynamics) are detrimental for coherent spin transfer.
In both GaAs and Si we have found a nonmonotonic dependence of $Q$ on $v$, with the lowest $Q$ at an optimal $v$. For GaAs in which the otherwise dominant dephasing due to hyperfine interaction with nuclei was weakened by monitoring of slow fluctuations of nuclear Overhauser field [51, 52, 53, 54], the most important shuttling-induced decoherence process is spin relaxation near the anticrossing. For Si, $Q$ sometimes exhibits two local minima, however, for most values of parameters considered here the global minimum of $Q$ occurs when processes of coherent and incoherent charge transfer error become comparably effective, i.e. the minimal coherent transfer error coincides with the minimal charge transfer error from [31]. A notable exception is the case of large $\Delta E_{Z}$, for which it is the competition between Landau-Zener process, and dephasing due to low-frequency noise in detuning that determines the optimal $v$ and minimal $Q$.
Applying our results to the case of shuttling for $\sim\!10$ $\mu$m distance (i.e. through a chain of $\sim\!100$ tunnel-coupled quantum dots), we have found that in silicon, coherent transfer with error $\leq\!10^{-3}$ will require $2t_{c}>60\,\mu\text{eV}$ for most dot pairs. When “weak links” with tunnel couplings twice smaller ($302t_{c}\!\approx\!$ 30 $\mu$eV) are present, qubit transport through them will dominate the total error and make it larger than $10^{-3}$. Shuttling through the “manipulation zones” near micromagnets is most error-prone, and weak links in these zones will be most detrimental for coherent shuttling. All these observations requirements for a hundred-dot array make the bucket brigade method of long-distance qubit transfer challenging for future designs, unless a method to obtain large and uniform tunnel coupling ($2t_{c}$ equal to at least 60 $\mu$eV), without having to tune up each tunnel barrier separately, is found.
@@ VI Summary
Acknowledgements.
This project was supported by the European Union’s Horizon 2020 research and innovation programme under grant agreement 101174557 (project QLSI2).
@@ Appendix B Relaxation rates
where $S_{i}(\Omega)$ is the spectral density of the environment, while the time-dependence of the relaxation rate $\Gamma_{sq\to s^{\prime}q^{\prime}}(t)$, that results from the time-dependence of the states |$\ket{\psi_{sq}(t)}$ and the energies $E_{sq}(t)$ was omitted for brevity.
@@
Using definition of the instantaneous states Eq. (A.1), we have:
@@ Appendix C Relaxation rates in Si and GaAs: Spectral densities
where we use $Z=50\Omega$ and $R_{q}=\pi/e^{2}\approx 13k\Omega$ [57].
@@
Figure 9: Spin relaxation rate in GaAs (a) and Si (b). In both panels we compare two tunnel couplings ($2t_{c}=40\mu$eV (solid line), $2t_{c}=100\mu$eV (dashed), as well as two spin splittings $E_{Z}=30\mu$eV (line without dots), $E_{Z}=10\mu$eV (lines with the dots). Finally for GaAs (a) we compare contributions from intrinsic SOC, i.e $|a|=1\mu$eV, $b=0$ (red) against synthetic $|a|=0,b_{\perp}=1\mu$eV (black). Similarly for Si (b) we compare intrinsic SOC $|a|=0.1\mu$eV, $b_{\perp}=0$ (green) against sythetic $|a|=0,b_{\perp}=1\mu$eV (yellow).
@@ Appendix D Dephasing due to low-frequency noise
We perform numerical test of an adiabatic dephasing. To do so we average $\varphi$, given by Eq. (86), over $\delta\epsilon$, $\delta t_{c}$, $\delta E_{Z,L}$, $\delta E_{Z,R}$. The results are depicted in Fig. 10, where numerically simulated $\langle\delta\varphi^{2}\rangle$ (dots) is shown as a function of noiseless difference in spin splittings $\Delta E_{Z}$ between the two dots.